OpenStudy (anonymous):

The surface area of a rectangular prism is found using the formula SA = 2(lw + lh + wh), where l is length, w is width, and h is height. What is the surface area (in square meters) of a rectangular prism with length 5 m, width 12 m, and height 6 m? A. 92 B. 102 C. 324 D. 360

OpenStudy (oaktree):

Just plug in.\[SA = 2(lw + lh + wh)\]Which is, in this case\[SA = 2(5*12 + 5*6 + 12*6)\]Which is the same as\[SA = 2(60 + 30 + 72)\]Which is\[SA = 2(162) = 324\]Which is answer choice C. Nothing that a calculator can't do.

OpenStudy (oaktree):

Thanks for the medal.

OpenStudy (anonymous):

thanks for the help

OpenStudy (oaktree):

No problem.

OpenStudy (anonymous):

The surface area of a rectangular prism is found using the formula SA = 2(lw + lh + wh), where l is length, w is width, and h is height. What is the surface area (in square meters) of a rectangular prism with length 5 m, width 12 m, and height 6 m?